# Efficient Routing of Inference Requests across LLM Instances in Cloud-Edge Computing

### v1

@@ Abstract
The rising demand for Large Language Model (LLM) inference services has intensified pressure on computational resources, resulting in latency and cost challenges. This paper introduces a novel routing algorithm based on the Non-dominated Sorting Genetic Algorithm II (NSGA-II) to distribute inference requests across heterogeneous LLM instances in a cloud-edge computing environment. Formulated as a multi-objective optimization problem, the algorithm balances response quality, response time, and inference cost, adapting to request heterogeneity (e.g., varying complexity and prompt lengths) and node diversity (e.g., edge vs. cloud resources). This adaptive routing algorithm optimizes performance under dynamic workloads. We benchmark the approach using a testbed with datasets including Stanford Question Answering Dataset (SQuAD), Mostly Basic Python Problems (MBPP), Hella Situations With Adversarial Generations (HellaSwag), and Grade School Math 8K (GSM8K). Experimental results show our solution, compared to the baselines, achieves up to 95.2% and 34.9% improvements in terms of response time and cost, respectively. These findings validate the algorithm’s effectiveness for scalable LLM deployments.
Index Terms:
large language model, inference request routing, meta-heuristic, cloud-edge computing
@@ IV-B3 Non-Dominated Sorting and Crowding Distance
NSGA-II employs non-dominated sorting to rank policies based on their dominance. A policy $\pi_{p}$ dominates $\pi_{q}$ if it is better in at least one objective and no worse in others. Policies are grouped into Pareto fronts, with the first front containing non-dominated solutions. To maintain diversity, NSGA-II computes the crowding distance for each policy within a front, measuring the Euclidean distance between neighboring solutions in the objective space ($RQ$, $C$, $RT$). Policies in less crowded regions are prioritized to ensure a well-distributed Pareto front [54].
@@ IV-B5 Algorithm Workflow
3:     Evaluate fitness of $P$ using $RQ$, $C$, and $RT$ (Equations (2), (3), (4))
@@ V-D Results and Analysis
[…]
@@
Based on the metrics and the resulting overall scores, we now examine how each routing strategy performs. From the data in Table II, the “Cloud Only” strategy achieves the highest average quality (0.5736) and lowest latency (1.0624 s) but also the highest cost (1.13 × $\times\,10^{-4}$), yielding an overall of 0.6667. In contrast, “Edge Only” minimizes cost (9.00 × $\times\,10^{-6}$) at the expense of quality (0.4207) and response time (3.9673 s), resulting in an overall of 0.3333. “Random Router” and “Round Robin Router” exhibit middling performance in all three dimensions, with overall scores of 0.3834 and 0.4062, respectively. Our “Proposed Router” attains average quality = 0.5462, average response_time = 1.1137 s, average cost = 7.36 × $\times\,10^{-5}$, which normalize to approximately $q_{\text{norm}}=0.8703$, $t_{\text{norm}}=0.9780$, $c_{\text{norm}}=0.4087$, yielding the highest overall of 0.7523. This confirms that our proposed router achieves the best aggregate trade‐off among quality, latency, and cost.
@@
  • Proposed Router: Inference quality $\approx 0.54$, response time $\approx 1.1\ \mathrm{s}$, cost $\approx 7\times 10^{-5}\ \$$. It lies almost as high and as far back as Cloud Only (indicating comparable quality and latency), yet much farther left (indicating significantly lower cost), thereby achieving the optimal overall trade‐off.
  • Cloud Only: Inference quality $\approx 0.58$ (highest) but cost $\approx 1.13\times 10^{-4}\ \$$ (highest). It appears at the “highest” and “farthest back” position, but is shifted rightward, indicating a higher cost.
  • Edge Only: Cost $\approx 9\times 10^{-6}\ \$$ (lowest), corresponding to the “farthest left” point; however, its inference quality $\approx 0.42$ (lowest) and response time $\approx 3.97\ \mathrm{s}$ (highest) place it at the “lowest” and “nearest” region, resulting in a weaker overall performance.
@@
In addition, we evaluate our proposed router’s performance under varying concurrency levels to assess scalability. The experiment begins with a concurrency equal to the number of nodes (4), then increasing to 8 and 10. Figure 4 illustrates the performance of our proposed router under concurrency levels 1, 4, 8, and 10. At a concurrency of 1, the router achieves baseline performance with a quality of 0.5462, response time of 1.1137 s, and a cost of 7.36 × $\times\,10^{-5}$. Increasing the concurrency to 4 slightly improves quality to 0.5536, with marginal increments in response time (1.1141 s) and negligible changes in cost (7.36 × $\times\,10^{-5}$), indicating effective load distribution at moderate concurrency. As concurrency rises further to 8, quality stabilizes at 0.5542, though response time (1.1660 s) and cost (7.40 × $\times\,10^{-5}$) begin to show moderate increases, suggesting emerging resource contention. At a concurrency level of 10, system capacity constraints become more pronounced, resulting in a reduced quality of 0.5438, elevated response time of 1.2061 s, and higher cost of 7.41 × $\times\,10^{-5}$. These results highlight the robustness and scalability of our proposed router under moderate concurrency conditions.
@@ VI Conclusion
Experimental evaluation on a realistic cloud-edge testbed using benchmarks such as SQuAD, MBPP, HellaSwag, and GSM8K demonstrated our method’s robustness and effectiveness. Specifically, our router achieved an average response quality of 0.5462, attaining 95.2% of the Cloud-Only baseline (0.5736), with only a modest 4.8% increase in response time (1.1137 vs. 1.0624 seconds). Significantly, inference costs were reduced by approximately 34.9% (from 1.13 × $\times\,10^{-4}$ to 7.36 × $\times\,10^{-5}$). Compared with baseline approaches such as Edge Only, Random Router, and Round Robin Router, our method excelled in efficiently balancing trade-offs among response quality, latency, and cost, maintaining consistent performance across varying concurrency scenarios.

### v2

@@ Abstract
The rising demand for Large Language Model (LLM) inference services has intensified pressure on computational resources, resulting in latency and cost challenges. This paper introduces a novel routing algorithm based on the Non-dominated Sorting Genetic Algorithm II (NSGA-II) to distribute inference requests across heterogeneous LLM instances in a cloud-edge computing environment. Formulated as a multi-objective optimization problem, the algorithm balances response quality, response time, and inference cost, adapting to request heterogeneity (e.g., varying complexity and prompt lengths) and node diversity (e.g., edge vs. cloud resources). This adaptive routing algorithm optimizes performance under dynamic workloads. We benchmark the approach using a testbed with datasets including Stanford Question Answering Dataset (SQuAD), Mostly Basic Python Problems (MBPP), Hella Situations With Adversarial Generations (HellaSwag), and Grade School Math 8K (GSM8K). Experimental results show our solution, compared to the baselines, preserves 95.2% of Cloud-Only response quality with slight latency increase, while reducing inference cost by 34.9%. These findings validate the algorithm’s effectiveness for scalable LLM deployments.
@@ IV-B3 Non-Dominated Sorting and Crowding Distance
NSGA-II employs non-dominated sorting to rank policies based on their dominance. A policy $\pi_{p}$ dominates $\pi_{q}$ if it is better in at least one objective and no worse in others. Policies are grouped into Pareto fronts, with the first front containing non-dominated solutions. To maintain diversity, NSGA-II computes the crowding distance for each policy within a front, measuring the Euclidean distance between neighboring solutions in the objective space ($RQ$, $C$, RTRT). Policies in less crowded regions are prioritized to ensure a well-distributed Pareto front [34].
@@ IV-B5 Algorithm Workflow
3:  Evaluate fitness of $P$ using RQRQ, $C$, and RTRT (Equations (2), (3), (4))
@@ V-D Results and Analysis
$Q_{i}$ $=\text{avg\_quality of router }i,$
$T_{i}$ $=\text{avg\_response\_time of router }i,$
$C_{i}$ $=\text{avg\_cost of router }i\,.$
@@
Based on the metrics and the resulting overall scores, we now examine how each routing strategy performs. From the data in Table II, the “Cloud Only” strategy achieves the highest average quality (0.5736) and lowest latency (1.0624 s) but also the highest cost (1.13 × $\times\,10^{-4}$ \$), yielding an overall of 0.6667. In contrast, “Edge Only” minimizes cost (9.00 × $\times\,10^{-6}$ \$) at the expense of quality (0.4207) and response time (3.9673 s), resulting in an overall of 0.3333. “Random Router” and “Round Robin Router” exhibit middling performance in all three dimensions, with overall scores of 0.3834 and 0.4062, respectively. Our “Proposed Router” attains average quality = 0.5462, average response_time = 1.1137 s, average cost = 7.36 × $\times\,10^{-5}$ \$, which normalize to approximately $q_{\text{norm}}=0.8703$, $t_{\text{norm}}=0.9780$, $c_{\text{norm}}=0.4087$, yielding the highest overall of 0.7523. This confirms that our proposed router achieves the best aggregate trade‐off among quality, latency, and cost.
@@
  • Proposed Router: Inference quality $\approx 0.54$, response time $\approx 1.1\ \mathrm{s}$, cost $\approx 7\times 10^{-5}\ \mathdollar$. It lies almost as high and as far back as Cloud Only (indicating comparable quality and latency), yet much farther left (indicating significantly lower cost), thereby achieving the optimal overall trade‐off.
  • Cloud Only: Inference quality $\approx 0.58$ (highest) but cost $\approx 1.13\times 10^{-4}\ \mathdollar$ (highest). It appears at the “highest” and “farthest back” position, but is shifted rightward, indicating a higher cost.
  • Edge Only: Cost $\approx 9\times 10^{-6}\ \mathdollar$ (lowest), corresponding to the “farthest left” point; however, its inference quality $\approx 0.42$ (lowest) and response time $\approx 3.97\ \mathrm{s}$ (highest) place it at the “lowest” and “nearest” region, resulting in a weaker overall performance.
@@
In addition, we evaluate our proposed router’s performance under varying concurrency levels to assess scalability. The experiment begins with a concurrency equal to the number of nodes (4), then increasing to 8 and 10. Figure 4 illustrates the performance of our proposed router under concurrency levels 1, 4, 8, and 10. At a concurrency of 1, the router achieves baseline performance with a quality of 0.5462, response time of 1.1137 s, and a cost of 7.36 × $\times\,10^{-5}$ \$. Increasing the concurrency to 4 slightly improves quality to 0.5536, with marginal increments in response time (1.1141 s) and negligible changes in cost (7.36 × $\times\,10^{-5}$ \$), indicating effective load distribution at moderate concurrency. As concurrency rises further to 8, quality stabilizes at 0.5542, though response time (1.1660 s) and cost (7.40 × $\times\,10^{-5}$ \$) begin to show moderate increases, suggesting emerging resource contention. At a concurrency level of 10, system capacity constraints become more pronounced, resulting in a reduced quality of 0.5438, elevated response time of 1.2061 s, and higher cost of 7.41 × $\times\,10^{-5}$ \$. These results highlight the robustness and scalability of our proposed router under moderate concurrency conditions.
@@ VI Conclusion
Experimental evaluation on a realistic cloud-edge testbed using benchmarks such as SQuAD, MBPP, HellaSwag, and GSM8K demonstrated our method’s robustness and effectiveness. Specifically, our router achieved an average response quality of 0.5462, attaining 95.2% of the Cloud-Only baseline (0.5736), with only a modest 4.8% increase in response time (1.1137 vs. 1.0624 seconds). Significantly, inference costs were reduced by approximately 34.9% (from 1.13 × $\times\,10^{-4}$ \$ to 7.36 × $\times\,10^{-5}$ \$). Compared with baseline approaches such as Edge Only, Random Router, and Round Robin Router, our method excelled in efficiently balancing trade-offs among response quality, latency, and cost, maintaining consistent performance across varying concurrency scenarios.